Monday, April 25, 2011

Nightlog 2011-04-24

SA: Alexei
SO: Siphelo
Others: Eric, Charl, Peter, Paul, Serafim


=> SALTICAM gripper has major problem. Does not communicate.
Only B-filter available for use.

We need someone from Cape Town to resolve this problem.

=> We were not able to make fine alignment and worked only with
coarse tonight. Charl, Peter and Paul were not able to help
finally. Hitesh promised by phone to check it tomorrow in remote

=> Clicky-click movements of the telescope did not work with SCAM.
We were need to make all movements by hands using "Tracking Adjust"

=> The focus shif between SALTICAM and RSS is -0.3mm following our
estimation with coarse stack.

=> Eric have amount of notes after his work with PCON, which he sent
to Ken and Anthony.


No comments:

Post a Comment